Ÿ=500+ 3xRegression analysis was applied between sales (Y in $1000) and advertising (X in $10,000), and the following estimated regression equation was obtained

Based on the above estimated regression line if advertising is $10,000, then the point estimated for sales (in dollars) is

a.503

b. 5030

c. 50,300

d.503,000
Answer #1

Since advertising spend is $10000, so putting x = 1, in the equation, we get,

\widehat{Y}=500+3(1)=503

i.e. 503 \times 1000=503000 i.e. $503000 since Y is in $10000

Hence option D.